OpenStudy (anonymous):

Leonard bought 9 CDs, bought 3 DVDs, and spent $15.43 on other items each DVD cost $14.99. Raj bought 15 CDs and had a coupon for 15% off his entire purchase. Both receipts had the same total. Write an equation and find out how much each CD costs.

OpenStudy (anonymous):

ill buy one of your cds

OpenStudy (anonymous):

what is it for?

OpenStudy (anonymous):

huh?

OpenStudy (austinl):

Okay, lets say that CDs are \(x\). Leonard bought 9 CDs, so \(9x\) he bought 3 DVDs at 14.99 each, so 44.97 for that, as well as 14.43 on miscellaneous items. That is equal to Raj's 15 CDs at a 15% discount. Do you think you could make an equation from this?

OpenStudy (austinl):

Make your equation, and show it to me. I will compare it to the one that I made and see if you are correct.

OpenStudy (anonymous):

I'm not sure what to do. :/

OpenStudy (austinl):

Here, \(9x+44.97+15.43=~?\) Do you think you could finish the equation?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

is it 44.97+15.43=60.4 9x=60.4 ? eh, I don't know. I'm not good at this. >.<

OpenStudy (anonymous):

I need help. :/

OpenStudy (austinl):

Raj had 15x at a discount of 15% \((15x)\times.15\) Now what would the entire equation look like?

OpenStudy (austinl):

Not quite.... \(9x+44.97+15.43=(15x).15\) Do you think you could solve for x?

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

How do I solve for x?